OpenStudy (anonymous):

If 10,500 is deposited in a compound interest account paying 2.99% interest annually, how much will be in the account after 5 years? Round your answer to the nearest cent.

jhonyy9 (jhonyy9):

10.500 ----- 100% x -----------2,99% --------------------- x = (2,99*10500)/100 = 313,95 this mony will get annually so what will be for 5 years 313,95*5=1569,75 - and you need assuming 10500 with 1569,75 for you get the right answer 10500 + 1569,75 = 12069,75 - so than will have after 5 years on account 12069,75 money

OpenStudy (anonymous):

Annual compound interest formula The formula for annual compound interest is A = P (1 + r/n) ^ nt: Where: A = the future value of the investment/loan, including interest P = the principal investment amount (the initial deposit or loan amount) r = the annual interest rate (decimal) n = the number of times that interest is compounded per year t = the number of years the money is invested or borrowed for

OpenStudy (anonymous):

Assuming it only compounds once per year $12,166.47 will be in the account after 5 years